Example 1: Using symmetry
X ~ N(50, 8²) and P(X < 58) = 0.8413. Find P(X > 58), P(X < 42) and P(42 < X < 58).
P(X > 58) = 1 − 0.8413 = 0.1587- 42 and 58 are both 8 from the mean, so by symmetry
P(X < 42) = P(X > 58) = 0.1587 P(42 < X < 58) = 1 − 2(0.1587) = 0.6826
Answer: 0.1587, 0.1587, 0.6826